# When Location Shapes Choice:

### v3

@@ 2 Model Formulation
Consider a ground set $N$ of $n$ substitutable products with fixed prices $(r_{i})_{i\in N}$. Let $\phi$ denote a choice model on the ground set $N$. We start with classic single-purchase discrete choice models where the customer chooses at most one product. The option of not selecting any product is symbolically represented as product 00, referred to as the no-purchase option. When any given assortment $S\subseteq N$ is offered to a customer, the choice model prescribes a probability of $\phi(i,S)$ for picking product $i\in S\cup\{0\}$ as the one to be purchased. Let $R(S)$ denote the expected revenue of an assortment $S\subseteq N$ which is given by
@@
We also consider multi-purchase choice models to capture settings where the customer may choose more than one product. For multi-purchase models, we use $\phi(i,S)$ to denote the probability that $i$ is one of the products purchased by the customer. Note that the expected revenue is still given by $R(S)=\sum_{i\in S}r_{i}\,\phi(i,S)$.
@@
Let $G=[m]$ denote a set of $m$ locations. Let $2^{G}$ denote the set of all subsets of locations. We have to place/display exactly one product at each location. As we discuss later, the constraint that there is at most one product at each location is without loss of generality (w.l.o.g.). We have many copies of each product and copies of a product can be placed/displayed at multiple locations in $G$. Let $X:{G}\to{N}$ denote a placement of products over $G$. More precisely, for each $j\in G$, the value $X(j)$ represents the product placed at location $j$. For notational convenience and to facilitate a more succinct representation, given a subset $L\subseteq G$, we use $X(L)$ to represent the set of (distinct) products placed at the locations specified by $L$, i.e., $X(L)=\cup_{j\in L}\{X(j)\}$.
@@
A customer randomly browses/visits a subset of locations on $G$ (see Figure 3 for an example) and then chooses from the products placed at the visited locations according to choice model $\phi$. Customers’ browsing behavior induces a distribution over $G$, we call this the browsing distribution $B$. Let $\mathbb{P}_{B}(L)$ denote the probability that the customer visits/browses the locations in $L$ in some order and then chooses from the set $X(L)\cup\{0\}$ of products (and no-purchase option). Let $j_{s}$ ($\geq 1$) denote the size of the smallest (non-empty) set with non-zero support in the browsing distribution. Similarly, let $j_{l}$ ($\leq m$) denote the size of the largest set with non-zero support in $\mathbb{P}_{B}$. Let $\rho=\frac{j_{l}+1}{j_{s}}$. This parameter captures the variation in the number of locations visited by different customers. Observe that $\rho\leq m+1$ and $\rho$ is independent of $m$ when $j_{l}$ is within a constant factor of $j_{s}$. W.l.o.g., let $\rho=\frac{j_{l}+1}{j_{s}}=2^{d}$ for some integer $d\in\{1,2,\cdots,\lceil\log_{2}(m+1)\rceil\}$. This assumption is without loss of generality because we can simply increase $j_{l}$ until $j_{l}=2^{d}\,j_{s}-1$, for some $d\geq 1$, by introducing dummy locations that are visited with zero probability.
@@
This is a standard assumption in the literature on assortment optimization for single-purchase choice models. The weak-rationality property indicates that the probability of choosing a specific product from a given assortment cannot increase when new products are added to the current assortment. Instead, it can only decrease, as people might potentially opt for the new product as a substitute. The monotonicity property states that in expectation, the number of purchased products increases as we add more products to the assortment. Assumption 2 is satisfied by a very general choice model known as the Random Utility model, which encompasses popular (single-purchase) choice models such as MNL, Markov Chain, Mixture of MNLs, Nested Logit. Assumption 2 is also satisfied by the multi-purchase choice models proposed by Bai et al. (2023) and Abdallah et al. (2024). Multi-purchase choice models that capture synergistic effects between complementary products do not satisfy Assumption 2 (Tulabandhula et al. (2023), Chen et al. (2022)). Our second assumption is as follows.
@@
Apart from the independence assumption above, we make no other structural assumptions on the probability distribution $B$, i.e., we allow for all possible customer browsing behavior over the set of $G$. Assumption 2 is also ubiquitous in the literature on special cases of Placement and other closely related problems (Davis et al. (2015), Gallego et al. (2020), Aouad and Segev (2021), Asadpour et al. (2023)). At the same time, there is also literature on models where the browsing distribution or the choice model is placement dependent (see Section 1.2).
@@
It will sometimes be simpler to define an algorithm without requiring that all locations are filled. Suppose we are given an infeasible placement solution where some locations are left empty. We describe a simple process to transform a placement with empty locations into a feasible assortment without decreasing the expected revenue of the placement. Let $i^{*}$ denote the product with the highest price, i.e.,
@@
From the above, the optimal solution to the relaxed problem is the same as that of the original problem. Thus, if $Y$ is a $\beta$-approximate solution to the relaxed problem, then $\hat{Y}$ is a $\beta$-approximate solution to the original Placement problem.
@@
From the hardness of product framing shown by Gallego et al. (2020), it follows that Placement on a Line is NP-hard. We include a proof in Appendix 7.
Lemma 2.4
Placement on a Line is weakly NP-hard even when the choice model $\phi$ is MNL.
In order to build our optimization framework for Placement, we assume access to two oracles. The first oracle computes an optimal or approximate solution for the cardinality constrained assortment optimization problem under the choice model $\phi$. The second oracle outputs any number of independent samples from the browsing distribution. In what follows, we formally present the two oracles.
@@ 2.1 Oracle for Cardinality Constrained Assortment Optimization
Our algorithms will use this oracle as a black box and the approximation guarantee of our algorithm will be proportional to $\alpha$. As we discussed in Section 1.2, there is an extensive literature cardinality constrained assortment optimization under various single-purchase choice models. Bai et al. (2023) introduce a multi-purchase choice model that satisfies Assumption 2 and give a PTAS for cardinality constrained assortment optimization. Abdallah et al. (2024) introduce a multi-purchase model based on the random utility framework and give a near-optimal algorithm for cardinality constrained assortment optimization in an asymptotic regime (large number of products).
@@ 2.2 Sampling Oracle for Browsing Distribution
For the sake of simplicity, we state and analyze the approximation guarantee of our algorithms assuming that we have exact knowledge of the expected revenue of any placement. In practice, the expectation may be hard to compute as the support of the browsing distribution can be as large as $2^{m}$, where $m$ is the number of locations. Our algorithms and their approximation guarantees hold even when we only have access to an oracle that outputs independent samples from the browsing distribution. We can use estimates of the expected revenue obtained by generating a small number (quantified later) of independent random samples of visited locations using a simulator (sampling oracle). In practice, one may use customer traffic data collected via video cameras or other sensors in physical stores and use browsing data in online stores to construct a simulator for the browsing distribution. We refer to Ziaei et al. (2023) and Flamand et al. (2023) for a detailed discussion on how this may be accomplished.
@@
Given parameters $(\epsilon,\delta)\in[0,1]^{2}$ and a placement $X$, the sample average revenue of $\frac{m^{2}\log\delta^{-1}}{2\epsilon^{2}}$ independent samples from the simulator gives an estimate ^$\hat{\mathcal{R}}(X)$ of the expected total revenue such that with probability $1-2\delta$,
^$\hat{\mathcal{R}}(X)\in[\mathcal{R}(X)-\epsilon\,\textsc{OPT},\mathcal{R}(X)+\epsilon\,\textsc{OPT}].$
@@
Given a set of placements $\{X_{1},X_{2},\cdots,X_{t}\}$, using $\frac{m^{2}\log\delta^{-1}}{2\epsilon^{2}}\cdot\log t$ samples from the simulator we can obtain revenue estimates {^$\{\hat{\mathcal{R}}(X_{1}),\cdots,\hat{\mathcal{R}}(X_{t})\}$ such that,
|^$|\hat{\mathcal{R}}(X_{k})-\mathcal{R}(X_{k})|\leq\epsilon\,\textsc{OPT}\quad\forall k\in[t],$
@@
Remarks on Usage of Exact Revenue Oracle: Recall that the approximation guarantee of our algorithms is stated under the assumption that we have exact knowledge of the revenue of any placement. Relaxing this assumption leads to a small loss in the approximation guarantee due to estimation errors. In particular, the final step in our main algorithm is to find the placement that maximizes $\hat{\mathcal{R}}(\cdot)$ from a small set of candidate placement solutions. From Lemma 2.7, given an approximation guarantee of $\beta$, our approximation guarantee in the absence of an exact revenue oracle is simply $\beta-2\epsilon$. In special cases such as when the product prices are identical (Section 3) or when the choice model is Markovian (Section 4.4), we propose greedy algorithms that construct a placement solution over several iterations. As we discuss later, the error due to estimates of revenue is sufficiently small even in these cases.
Deterministic vs. Randomized Algorithms: Given an exact revenue oracle, we say that an algorithm is deterministic if it does not use any randomness. In contrast, a randomized algorithm uses additional randomness and may output different solutions each time the algorithm is executed.
@@ 3 Uniformly Priced Products
In this section, we focus on single-purchase choice models that belong to the general class of Random Utility Models (RUM) and we tackle the Placement problem with uniformly priced products. In particular, we present a $(1-1/e)$ approximation algorithm for the problem in Section 3.1. In Section 3.2, we show that this approximation is tight; that is, the Placement problem with uniformly priced products cannot be approximated within a factor better than $(1-\frac{1}{e}+\epsilon)$ for any $\epsilon>0$, unless P=NP. The hardness result follows from the hardness of cardinality constrained assortment optimization under a (single-purchase) Mixture of MNL (MMNL) model when products have uniform prices.
@@ 3.1 Placement of Uniformly Priced Products
For any RUM choice model, the expected revenue function $R(\cdot)$ is both monotone and submodular when all products are priced identically.
Lemma 3.1 is a well-established result observed across various RUM choice models. As an example, Désir et al. (2020) demonstrates this result for the Markov Chain choice model. Berbeglia and Joret (2020) provides a proof of this lemma for general RUM choice models (see Lemma 4.2 in their paper). Note that there are choice models that are not RUM but adhere to Assumption 2, in which the revenue function is not submodular, and therefore do not verify the above lemma. Berbeglia and Joret (2020) presents such an example in their paper. Hence, we restrict ourselves to RUM models in this section.
Building on Lemma 3.1, we establish that the Placement problem with identical prices, can be formulated as a problem of maximizing a monotone submodular function under a matroid constraint [^1] A matroid on ground set $U$ is given by a family of subsets of $U$, denoted $\mathcal{F}$. Every set in $\mathcal{F}$ is called an independent set of the matroid. $\mathcal{F}$ satisfies the following properties: (i) The empty set, $\phi$, is independent. (ii) If $A\in\mathcal{F}$ then $B\in\mathcal{F}$ for all $B\subseteq A$. (iii) If $A,B\in\mathcal{F}$ and $|B|<|A|$, then $B\cup\{x\}\in\mathcal{F}$ for some $x\in A\backslash B$.. There are several algorithms for maximizing a monotone submodular function subject to a matroid constraint. Nemhauser et al. (1978) showed that a greedy algorithm that iteratively builds the solution by adding elements one by one is $0.5$ approximate. Calinescu et al. (2011) give $(1-1/e)$-approximation algorithms for the problem which is the best known approximation. We refer to Badanidiyuru and Vondrák (2014) for a review of other algorithms for the problem. The next theorem shows how to obtain $(1-1/e)$-approximation algorithm for the Placement problem with identical prices.
@@
There is a $(1-1/e)$-approximation algorithm for the Placement problem with uniformly priced products and a RUM choice model.
We include the proof of Theorem 3.2 in Appendix 8. Note that Asadpour et al. (2023) show a $(1-1/e)$ guarantee for a similar problem setting by generalizing ideas from submodular optimization. The browsing distribution in their setting is the same as Placement on a Line but they consider additional constraints arising out of fairness considerations. Our result is for a general browsing distribution without the latter constraints.
@@
There is a $\frac{O(1)}{\log\frac{r_{\max}}{r_{\min}}}$ approximation algorithm for Placement with arbitrary prices under a RUM choice model.

### v4

@@ 2 Model Formulation
Consider a ground set $N$ of $n$ substitutable products with fixed prices $(r_{i})_{i\in N}$. Let $\phi$ denote a choice model on the ground set $N$. We start with classic single-purchase discrete choice models where the customer chooses at most one product. The option of not selecting any product is symbolically represented as product 0, referred to as the no-purchase option. When any given assortment $S\subseteq N$ is offered to a customer, the choice model prescribes a probability of $\phi(i,S)$ for picking product $i\in S\cup\{0\}$ as the one to be purchased. Let $R(S)$ denote the expected revenue of an assortment $S\subseteq N$ which is given by
@@
We also consider multi-purchase choice models to capture settings where the customer may choose more than one product. For multi-purchase models, we use $\phi(i,S)$ to denote the expected number of times that the customer purchases product $i$. Note that the expected revenue is still given by $R(S)=\sum_{i\in S}r_{i}\,\phi(i,S)$.
@@
Let $G=[m]$ denote a set of $m$ locations. Let $2^{G}$ denote the set of all subsets of locations. We have to place/display at most one product at each location. As we discuss later (Remarks 2.2 and 2.3), this constraint is without loss of generality (w.l.o.g.) in our formulation. We have many copies of each product and copies of a product can be placed/displayed at multiple locations in $G$. Let $X:{G}\to{N}$ denote a placement of products over $G$. More precisely, for each $j\in G$, the value $X(j)$ represents the product placed at location $j$. For notational convenience and to facilitate a more succinct representation, given a subset $L\subseteq G$, we use $X(L)$ to represent the set of (distinct) products placed at the locations specified by $L$, i.e., $X(L)=\cup_{j\in L}\{X(j)\}$. configuration.
@@
A customer randomly browses/visits a subset of locations on $G$ (see Figure 3 for an example) and then chooses from the products placed at the visited locations according to choice model $\phi$. Customers’ browsing behavior induces a distribution over $G$, we call this the browsing distribution $B$. Let $\mathbb{P}_{B}(L)$ denote the probability that the customer visits/browses the locations in $L$ in some order and then chooses from the set $X(L)\cup\{0\}$ of products (and no-purchase option). Let $j_{\min}$ ($\geq 1$) denote the size of the smallest (non-empty) set with non-zero support in the browsing distribution. Similarly, let $j_{\max}$ ($\leq m$) denote the size of the largest set with non-zero support in $\mathbb{P}_{B}$. Let $\rho=\frac{j_{\max}+1}{j_{\min}}$. This parameter captures the variation in the number of locations visited by different customers. Observe that $\rho\leq m+1$ and $\rho$ is independent of $m$ when $j_{\max}$ is within a constant factor of $j_{\min}$. W.l.o.g., let $\rho=\frac{j_{\max}+1}{j_{\min}}=2^{d}$ for some integer $d\in\{1,2,\cdots,\lceil\log_{2}(m+1)\rceil\}$. This assumption is without loss of generality because we can simply increase $j_{\max}$ until $j_{\max}=2^{d}\,j_{\min}-1$, for some $d\geq 1$, by introducing dummy locations that are visited with zero probability.
@@
This is a standard assumption in the literature on assortment optimization for single-purchase choice models. The weak-rationality property indicates that the probability of choosing a specific product from a given assortment cannot increase when new products are added to the current assortment. Instead, it can only decrease, as people might potentially opt for the new product as a substitute. The monotonicity property states that in expectation, the number of purchased products increases as we add more products to the assortment. Assumption 2 is satisfied by all choice models that fall within the RUM framework, which encompasses popular (single-purchase) choice models such as MNL, Markov Chain, Mixture of MNLs, Nested Logit. Assumption 2 is also satisfied by the multi-purchase choice models proposed by Bai et al. (2024) and Abdallah et al. (2024). Multi-purchase choice models that capture synergistic effects between complementary products do not satisfy Assumption 2 (Tulabandhula et al. (2023), Chen et al. (2022), El Housni et al. (2025)). Our second assumption is as follows.
@@
Apart from the independence assumption above, we make no other structural assumptions on the probability distribution $B$, i.e., we allow for all possible customer browsing behavior over the set of $G$. Assumption 2 is also ubiquitous in the literature on special cases of Placement and other closely related problems discussed in Section 1.2 (Davis et al. (2015), Gallego et al. (2020), Aouad and Segev (2021), Asadpour et al. (2023)). At the same time, there is also literature on models where the browsing distribution or the choice model is placement dependent (see Section 1.2).
@@
In order to build our optimization framework for Placement, we assume access to two oracles. The first oracle computes an optimal or approximate solution for the cardinality constrained assortment optimization problem under the choice model $\phi$. The second oracle outputs any number of independent samples from the browsing distribution. In the following sections, we formally present the two oracles.
Remark 2.1 (Product Repetition)
Placing a product in multiple locations is common in both physical and digital retail environments. In grocery stores, for instance, a product might appear on end-cap displays, checkout lanes, or promotional aisles. Big-box retailers often use promotional zones or pallet displays to emphasize high-volume or discounted items. Apparel stores highlight featured styles using mannequins and dedicated displays. Similarly, e-commerce platforms showcase repeated products in “best-selling” or “featured” sections, often at the top of the page. As illustrated in Figure 2, a single product search may return multiple listings for the same base item, differentiated by size, color, or configuration.
@@
Since we are allowed to repeat a product at multiple locations, we can ensure that all locations are filled without losing revenue. Suppose we are given a placement solution where some locations are left empty. We describe a simple process to transform a placement with empty locations into a feasible assortment without decreasing the expected revenue of the placement. Let $i^{*}$ denote the product with the highest price, i.e.,
@@
From the above, the optimal solution to Placement has the same revenue as the optimal solution to Placement with the additional constraint that all locations must be filled. Thus, if $Y$ is a $\beta$-approximate solution to Placement then $\hat{Y}$ is a $\beta$-approximate solution to the constrained problem as well.
@@
For Placement on a Line, there is no advantage in repeating a product: If a product is at location $j$, then a copy of the product at location $j+k$ does not contribute additional revenue because a customer who visits location $j+k$ will also visit location $j$. So if a product is repeated at $j$ and $j+k$ then we can simply leave $j+k$ empty without changing revenue. In Section 4.2, we show that repeating products makes a big difference to the revenue for the general Placement problem.
@@ 2.1 Oracle for Cardinality Constrained Assortment Optimization
Our algorithms will use this oracle as a black box and the approximation guarantee of our algorithm will be proportional to $\alpha$. As we discussed in Section 1.2, there is an extensive literature cardinality constrained assortment optimization under various single-purchase choice models. Bai et al. (2024) introduce a multi-purchase choice model that satisfies Assumption 2 and give a PTAS for cardinality constrained assortment optimization. Abdallah et al. (2024) introduce a multi-purchase model based on the RUM framework and give a near-optimal algorithm for cardinality constrained assortment optimization in an asymptotic regime (large number of products).
@@ 2.2 Sampling Oracle for Browsing Distribution
For the sake of simplicity, we state and analyze the approximation guarantee of our algorithms assuming that we have exact knowledge of the expected revenue of any placement. In practice, the expectation may be hard to compute as the support of the browsing distribution can be as large as $2^{m}$, where $m$ is the number of locations. Our algorithms and their approximation guarantees hold even when we only have access to an oracle that outputs independent samples from the browsing distribution. We can use estimates of the expected revenue obtained by generating a small number (quantified later) of independent random samples of visited locations using a sampling oracle for the browsing distribution.
@@
Given parameters $(\epsilon,\delta)\in[0,1]^{2}$ and a placement $X$, the sample average revenue of $\frac{m^{2}\log\delta^{-1}}{2\epsilon^{2}}$ independent samples from the sampling oracle gives an estimate $\hat{\mathcal{R}}(X)$ of the expected total revenue such that with probability $1-2\delta$,
$\hat{\mathcal{R}}(X)\in[\mathcal{R}(X)-\epsilon\,\textsc{OPT},\mathcal{R}(X)+\epsilon\,\textsc{OPT}].$
@@
Given a set of placements $\{X_{1},X_{2},\cdots,X_{t}\}$, using $\frac{m^{2}\log\delta^{-1}}{2\epsilon^{2}}\cdot\log t$ samples from the sampling oracle we can obtain revenue estimates $\{\hat{\mathcal{R}}(X_{1}),\cdots,\hat{\mathcal{R}}(X_{t})\}$ such that,
$|\hat{\mathcal{R}}(X_{k})-\mathcal{R}(X_{k})|\leq\epsilon\,\textsc{OPT}\quad\forall k\in[t],$
@@
Remark 2.8 (Usage of Exact Revenue Oracle)
[…]
From Lemma 2.7, given an approximation guarantee of $\beta$, our approximation guarantee in the absence of an exact revenue oracle is simply $\beta-2\epsilon$. In special cases such as when the product prices are identical (Section 3) or when the choice model is Markovian (Section 4.4), we propose greedy algorithms that construct a placement solution over several iterations. As we discuss later, the error due to estimates of revenue is sufficiently small even in these cases.
Remark 2.9 (Deterministic vs. Randomized Algorithms)
Given an exact revenue oracle, we say that an algorithm is deterministic if it does not use any randomness. In contrast, a randomized algorithm uses additional randomness and may output different solutions each time the algorithm is executed.
Remark 2.10 (Constructing the Sampling Oracle from Data)
In practice, one may use customer traffic data collected via RFID (Larson et al. 2005) and other technologies (Ziaei et al. 2023) in physical stores, and browsing data in online stores, to construct a sampling oracle for the browsing distribution. We refer to Larson et al. (2005) and Ziaei et al. (2023) for more details on how these technologies are used in physical stores. Flamand et al. (2023) propose a different approach based on learning a parametric model of customers’ browsing behavior using sales transaction data. The idea of learning something about the customer’s path in the store from their set of purchased products is based on the assumption that customers shop in a way that minimizes their walking distance (Hui et al. 2009). Under this assumption, given the store layout, the product placement, and set of products purchased by a customer, one can use optimization techniques to infer the customer’s path through the store. Flamand et al. (2023) use this technique to learn the parameters of a model that predicts the probabilities of a customer visiting a product in a given location. For a fixed product placement our model reduces to a consider-then-choose choice model and one can estimate the revenue by using techniques for estimating consider-then-choose models (Jagabathula et al. 2024).
@@ 3 Uniformly Priced Products
In this section, we focus on single-purchase choice models that belong to the general framework of Random Utility Maximization (RUM) and we tackle the Placement problem with uniformly priced products. In particular, we present a (11/e)(1-1/e) approximation algorithm for the problem in Section 3.1. In Section 3.2, we show that this approximation is tight; that is, the Placement problem with uniformly priced products cannot be approximated within a factor better than $(1-\frac{1}{e}+\epsilon)$ for any $\epsilon>0$, unless P=NP. The hardness result follows from the hardness of cardinality constrained assortment optimization under a (single-purchase) Mixture of MNL (MMNL) model when products have uniform prices.
@@ 3.1 Placement of Uniformly Priced Products
For any RUM-based choice model, the expected revenue function $R(\cdot)$ is both monotone and submodular when all products are priced identically.
Lemma 3.1 is a well-established result observed across various RUM-based choice models. As an example, Désir et al. (2020) demonstrates this result for the Markov Chain choice model. Berbeglia and Joret (2020) provides a proof of this lemma for general RUM-based choice models (see Lemma 4.2 in their paper). Note that there are choice models that are not RUM-based but adhere to Assumption 2, in which the revenue function is not submodular, and therefore do not verify the above lemma. Berbeglia and Joret (2020) presents such an example in their paper. Hence, we restrict ourselves to RUM-based choice models in this section.
Building on Lemma 3.1, we establish that the Placement problem with identical prices, can be formulated as a problem of maximizing a monotone submodular function under a matroid constraint [^1] A matroid on ground set $U$ is given by a family of subsets of $U$, denoted $\mathcal{F}$. Every set in $\mathcal{F}$ is called an independent set of the matroid. $\mathcal{F}$ satisfies the following properties: (i) The empty set, $\phi$, is independent. (ii) If $A\in\mathcal{F}$ then $B\in\mathcal{F}$ for all $B\subseteq A$. (iii) If $A,B\in\mathcal{F}$ and $|B|<|A|$, then $B\cup\{x\}\in\mathcal{F}$ for some $x\in A\backslash B$.. There are several algorithms for maximizing a monotone submodular function subject to a matroid constraint. Nemhauser et al. (1978) showed that a greedy algorithm that iteratively builds the solution by adding elements one by one is $0.5$ approximate. Calinescu et al. (2011) give (11/e)(1-1/e)-approximation algorithms for the problem which is the best known approximation. We refer to Badanidiyuru and Vondrák (2014) for a review of other algorithms for the problem. The next theorem shows how to obtain (11/e)(1-1/e)-approximation algorithm for the Placement problem with identical prices.
@@
There is a (11/e)(1-1/e)-approximation algorithm for the Placement problem with uniformly priced products and a RUM-based choice model.
We include the proof of Theorem 3.2 in Appendix 8. Note that Asadpour et al. (2023) show a (11/e)(1-1/e) guarantee for a similar problem setting by generalizing ideas from submodular optimization. The browsing distribution in their setting is the same as Placement on a Line but they consider additional constraints arising out of fairness considerations. Our result is for a general browsing distribution without the latter constraints.
@@
There is a $\frac{O(1)}{\log\frac{r_{\max}}{r_{\min}}}$ approximation algorithm for Placement with arbitrary prices under a RUM-based choice model.